OpenStudy (anonymous):

2 sides of an equilateral triangle have lenghts of 3x+1 and -2x+26. could 21- x or 3x-4 be the lenght of the 3rd side? a. both 21-x and 3x-4 b. 21-x only c. neither 21-x nor 3x-4 d. 3x-4 only

ganeshie8 (ganeshie8):

hint : in equilateral triangle, all sides are equal

ganeshie8 (ganeshie8):

they gave u 2 sides.... set them equal, and solve x first

OpenStudy (anonymous):

3x+1+-2x+26??

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yu need to set them equal, cuz sides are equal

ganeshie8 (ganeshie8):

like this : 3x+1 = -2x+26

ganeshie8 (ganeshie8):

solve x now

OpenStudy (anonymous):

oh okay, 3x+1=-2x+26 +x +x

OpenStudy (anonymous):

im sorry im a little slow

OpenStudy (anonymous):

x=5

ganeshie8 (ganeshie8):

it's :)

ganeshie8 (ganeshie8):

3x+1=-2x+26

ganeshie8 (ganeshie8):

add 2x both sides

ganeshie8 (ganeshie8):

3x+1=-2x+26 +2x +2x

ganeshie8 (ganeshie8):

5x + 1 = 26

ganeshie8 (ganeshie8):

subtract 1

OpenStudy (anonymous):

substitute the value of x, =16 both sides :)

OpenStudy (anonymous):

5x+1=-4x+26 -1 -1

ganeshie8 (ganeshie8):

yes, keep going

ganeshie8 (ganeshie8):

hey no wait

ganeshie8 (ganeshie8):

we got this right, 5x + 1 = 26

ganeshie8 (ganeshie8):

subtract 1

ganeshie8 (ganeshie8):

5x + 1 = 26 -1 -1 5x = 25

OpenStudy (anonymous):

-2x+26=3x+1 -2x-3x=1-26 -5x=-25 x=5 then substitute the value of x

OpenStudy (anonymous):

side 1 3x+1= 3(5)+1 =16 side 2 -2x+26= -2(5)+26 =-10+26 =16

OpenStudy (anonymous):

so its a

OpenStudy (anonymous):

@engr04

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@terenzreignz

ganeshie8 (ganeshie8):

read the q once, could 21- x or 3x-4 be the lenght of the 3rd side?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

it's b

ganeshie8 (ganeshie8):

you got x = 5

OpenStudy (anonymous):

21-x 21-5=16 correct but, 3x-4 3(5)-4=11 incorrect so it's only 21-x the answer is b :)